Circular No. 7242

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

NOVA CIRCINI 1999 W. Liller, Viña del Mar, Chile, reports his discovery of an apparent

nova (m_{pv} = 7.7) on Technical Pan photographs taken with an 85-mm camera lens (+ orange filter; via PROBLICOM) on Aug. 23.0110 and 23.0125 UT. CCD frames taken with a 0.20-m Schmidt camera on Aug. 23.9717 yield a position of α = 14^{h}23^{m}23^{s}.6, δ = −69^{o}08′45′′.1 (equinox 2000.0; uncertainty ± 0′′.5), at which time the magnitude measured with a broadband V filter was 7.79. Nothing brighter than mag 11.0 appears at this location on two photographs taken on July 15.99. A low-resolution spectrogram obtained with the Schmidt (+ objective prism + CCD) on Aug. 23.9874 shows Hα in emission at a level ∼ 0.85 times brighter than the intensity of the surrounding continuum. The modest strength of this emission suggests that the nova outburst occurred only a day or two earlier. A. Pearce, Nedlands, W. Australia, reports the star at visual mag 8.1 on Aug. 24.581, noting that nothing is visible on the Digital Sky Survey field at the position of the nova. Other visual magnitude estimates: Aug. 24.037, 7.5 (R. Stubbings, Drouin, Victoria, Australia); 24.302, 8.4 (A. Jones, Nelson, New Zealand).

SUPERNOVA 1999do IN MARKARIAN 922 M. Modjaz and W. D. Li, University of California at Berkeley, on

behalf of the Lick Observatory Supernova Search (cf. IAUC 6627, 7126), report the discovery with the 0.8-m Katzman Automatic Imaging Telescope (KAIT) of an apparent supernova in Mrk 922 (= MCG +05-54-003) on unfiltered images taken on Aug. 20.3 (mag ∼ 18.0) and 21.3 UT (mag ∼ 17.4). The new object is located at α = 22^{h}53^{m}29^{s}.87, δ = +31^{o}38′42′′.9 (equinox 2000.0), which is 4′′.7 east of the nucleus of Mrk 922. A KAIT image of the same field taken on Aug. 16.3 (limiting mag ∼ 19.5) showed nothing at this position.

COMET 10P/TEMPEL 2 Visual m_{1} estimates: Aug. 2.08 UT, 10.6 (J. Bortle, Stormville, NY,

0.41-m reflector); 9.36, 9.5 (S. T. Rae, Palmerston North, New Zealand, 0.25-m reflector); 13.85, 8.8 (M. Reszelski, Szamotuly, Poland, 20×60 binoculars); 22.56, 9.2 (A. Pearce, Nedlands, W. Australia, 20×80 binoculars).

1999 August 24  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green